OpenStudy (abbycross167):

Can someone please help me on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

State the domain of y = x^2

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you can square any number, so unless you make your own restriction the domain would be \[\mathbb{R}\]

OpenStudy (abbycross167):

the domain would just be R? no numbers?

OpenStudy (abbycross167):

@mathmate could you help me?

OpenStudy (fifciol):

yes, the domain is simply R

OpenStudy (abbycross167):

So for the answer I just put R? No steps or anything?

OpenStudy (abbycross167):

@Fifciol

OpenStudy (fifciol):

I think so

OpenStudy (mathmate):

R means all real numbers. Yes, the domain of y=x^2 is all real numbers, usually written as \(]-\infty,+\infty[\), or simply R. Note that all polynomials with real coefficients have \(]-\infty,+\infty[\), or R as domain. Since x^2 is a 2nd degree polynomial with one single term and real coefficient, the domain is no exception, i.e. R.